# Renovating Names in Open-Vocabulary Segmentation Benchmarks

### v1

@@ Abstract
Names are essential to both human cognition and vision-language models. Open-vocabulary models utilize class names as text prompts to generalize to categories unseen during training. However, name qualities are often overlooked and lack sufficient precision in existing datasets. In this paper, we address this underexplored problem by presenting a framework for “renovating” names in open-vocabulary segmentation benchmarks (RENOVATE). Through human study, we demonstrate that the names generated by our model are more precise descriptions of the visual segments and hence enhance the quality of existing datasets by means of simple renaming. We further demonstrate that using our renovated names enables training of stronger open-vocabulary segmentation models. Using open-vocabulary segmentation for name quality evaluation, we show that our renovated names lead to up to 16% relative improvement from the original names on various benchmarks across various state-of-the-art models. We provide our code and relabelings for several popular segmentation datasets (ADE20K, Cityscapes, PASCAL Context) to the research community.
Keywords:
Vision-Language Datasets Open-Vocabulary Segmentation Renaming
@@ 1 Introduction
Figure 1: Problems of names in current segmentation benchmarks. We demonstrate with examples from well-known datasets: MS COCO [22], ADE20K [57], and Cityscapes [7]. Our renovated names are more aligned with the visuals and help models to generalize better.
In this work, we focus on renovating the names for open-vocabulary segmentation benchmarks, as their dense annotations pose a greater challenge and offer broader applicability than other recognition tasks. The importance of names in open-vocabulary segmentation is often overlooked, with one exception, OpenSeg [11], which manually inspected and modified the class names of several segmentation benchmarks, resulting in considerable performance gains of their proposed model. While their work demonstrated the importance of choosing good names, a manual approach is subjective and difficult to scale to finer granularity. To date, there exists no systematic study on how to rename benchmarks in a scalable and principled way. Our work is the first attempt to tackle this challenge.
In this work, we introduce a scalable, simple, yet general method for renaming segmentation benchmarks. Our approach leverages foundation models to automate the renaming process, significantly reducing the manual labor traditionally required. Towards this goal, our method can be divided into two steps. First, we narrow down the name search space from the whole language space to a curated list of candidate names. This list is generated by leveraging the original class names with contextually relevant nouns extracted from visual contents by an image captioning model, thereby streamlining the search for names. Next, we employ a specially trained renaming model to identify and select the best-matching candidate name for every ground-truth segmentation mask. In this way, we match an individual name for each instance without any extra human annotations.
To validate the quality of our renovated names, we first conduct a careful human preference study showing that our method successfully finds more accurate names for current segmentation benchmarks. We then apply our renovated names to upgrade existing benchmarks, resulting in datasets that not only contain more precise names of the visual segments, but also expand the number of classes by $4\sim 5$ times. This enhancement makes the benchmarks significantly more challenging and realistic than the original benchmarks. Then, we use our renovated names for training open-vocabulary models and show that our names improve the open-vocabulary generalization performances on various benchmarks. Finally, we leverage pre-trained open-vocabulary models to automate the name quality evaluation process. Our results on the ADE20K, Cityscapes, and PASCAL Context benchmarks show that the renovated names improve pre-trained models up to 16% over the original names, suggesting that our names match more accurately with the visual segments.
@@
  • We point out the name problems in existing open-vocabulary segmentation benchmarks.
  • We propose a scalable, simple, general framework (RENOVATE) for renaming segmentation benchmarks.
  • We demonstrate that our renovated names are of high quality and can upgrade existing benchmarks to be more challenging and realistic, as well as help train models with stronger open-vocabulary capabilities.
@@ 3.1 Generating candidate names
We use GPT-4 [28] for creating a pool of candidate names. A naive solution is to prompt GPT-4 with the original name and ask it to generate synonyms and hierarchical concepts. However, since the original names are often too general and ambiguous, GPT-4 does not have sufficient knowledge to generate high-quality candidates. Therefore, we propose to exploit the visual contents for generating some “context names” (Tab. 1) that assist GPT-4 in comprehending the category’s meaning prior to generating candidate names.
@@
Specifically, we instruct GPT-4 to focus on two types of candidate name generation: (1) Synonyms or subcategories, which can expand the lexical scope of the overly general class name. For instance, the class name “grass” could yield variations like “lawn” or “turf”. For this type, synonyms and subcategories are considered together, acknowledging their often interchangeable nature in certain contexts. (2) Providing short, distinct contexts to polysemous class names. As an example, the ambiguous class name “fan” could be elaborated into “ceiling fan” or “floor fan”, offering clearer specificity. As shown in Tab. 1, the 5-10 candidate names generated by GPT-4 are more precise than the original names, reflecting the context information encoded by the context names. We provide detailed GPT-4 prompts in the supplement.
@@ 3.2 Training for candidate name selection
When using GPT-4 to generate candidate names, generating a surplus is not an issue. Among the generated candidates, we will only keep those that are better aligned with the visual contents from the images. Particularly, we train a model that can assess the alignment between each name and segment. For training such a model, we use segmentation as the proxy task. Conceptually, a candidate name that well describes the segment in the image should help the segmentation model recover its ground-truth mask and make a correct classification. Therefore, the model should allow vision-language interaction, and the supervision signal should encourage to use textual information for mask prediction and classification.
Our renaming model is illustrated in Fig. 2. Its vision part starts from a CLIP-based vision backbone followed by a pixel decoder that gradually upscales the backbone features to generate per-pixel embeddings $F_{\text{pix}}\in\mathbb{R}^{H\times W\times C}$. The candidate names of the categories present in the image are processed by the CLIP text encoder. The interaction of textual and visual information takes place at the transformer decoder. The initial queries $X^{(0)}$ are from the text embeddings, and then updated by the transformer decoder consisting of $L$ blocks using the multi-scale visual features from the pixel decoder. At the output of each block, the queries $X^{(l)}$ generate intermediate mask and class predictions as follows:
@@
where *** refers to pixel-wise multiplication. The final predictions are made by the queries $X^{(L)}$. Our pixel decoder, mask prediction and classification follow the design of Mask2Former [4], which handles panoptic, instance, and semantic segmentation in a unified manner. The key differences lie in 1) the transformer decoder design, and 2) the training strategy, which we detail subsequently.
Transformer Decoder. Our transformer decoder uses text embeddings as input queries, unlike most open-vocabulary segmentation models that employ text embeddings solely for mask classification as weights of the Linear layer in Eq. 1. This key change fosters an earlier integration of text and visual features at varying scales, essential for leveraging text information in segmentation tasks. As shown in Fig. 2, the text embedding of each candidate name from the segment’s class “bicycle” feeds into the transformer decoder, generating corresponding mask and class predictions for each candidate name.
@@
Training Strategy. With both the vision and text encoder of CLIP kept frozen, we train the transformer decoder together with the pixel decoder for segmentation. To recover the ground-truth mask $M_{i}$ and class $c_{i}$, the transformer decoder makes multiple mask and class predictions with candidate names from the class $c_{i}$. The prediction with the highest Intersection over Union (IoU) with $M_{i}$ is selected for loss computation:
@@ 3.3 Obtaining renovated names
As illustrated in Fig. 2(a), to obtain the renovated names, we run the trained renaming model to associate each ground-truth mask with the candidate name. For each segment, our renaming model first ranks the names by the IoU scores between the predicted masks and the ground-truth mask, then selects the top-scoring name as the renovated name. While the IoU ranking scores are not probabilities, they are valuable to understand the relative and absolute model confidence in the names, i.e., how the chosen name stands against the other candidates and whether the selection is indeed confident. In the experiments (Sec. 4.2), we make use of this information to speed up human verification. We show some examples of renovated names in Fig. 4 and more in the supplements.
@@
Our renovated names are readily usable for analyzing the dataset in a more fine-grained manner, as exemplified in Fig. 2(b). This can be helpful for countering the dataset biases during dataset curation and collection. Note that our renovated names in each original class are not necessarily mutually exclusive. In fact, we intentionally did not include the mutual exclusivity as a constraint during candidate name generation because different visual segments may be matched to names at different hierarchies which are not mutually exclusive. For example, in Fig. 4, a person is named as “engineer” (second row), likely due to the clothing, while the other person is named as “woman” (third row), possibly from the physical appearance. But “engineer” and “woman” are not mutually exclusive names. In the future, it would be an interesting extension to further study the hierarchies and other relationships of different names in a dataset.
@@ 4 Experiments
Datasets. We apply RENOVATE to MS COCO [22], ADE20K [57], Cityscapes [7], and PASCAL Context [27]. Both MS COCO and Cityscapes have panoptic annotations. ADE20K has two annotation sets: ADE-150 with panoptic annotations for 150 classes and ADE-847 with only semantic annotations for 847 classes. PASCAL Context has only semantic annotations for 459 classes. When the panoptic annotations are available, we perform RENOVATE on panoptic segments, otherwise, we rename the semantic segments.
@@
Implementation details. To generate context names, our default image captioning model is CaSED [6], which retrieves captions by matching vision embeddings with text embeddings of captions from a large-scale external image-caption dataset PMD [39]. For each dataset to be renamed, we train a renaming model for 60k iterations with a batch size of 16 only on the training set. See details in the supplements.
Name baselines. We consider the only two other existing name baselines for open-vocabulary segmentation benchmarks: (1) Original class names; (2) OpenSeg class names, which were manually curated by previous researchers [11].
(a) User interface for human preference study.
(b) Human preference results.
Figure 5: Human preference study. A survey of 20 researchers is conducted to compare preferences between the original names versus RENOVATE names on the validation sets. RENOVATE names are favored in 82% cases.
4.1 Human preference study
As demonstrated in Fig. 5, we first study the name quality of our renovated names with a human preference study. Specifically, we first randomly select 10% of all the segments in the validation set of each dataset. Then for each segment, we ask 20 researchers unfamiliar with the project about their preference between its original name and our model-selected name. Our results demonstrate that our renovated names are preferred in 82% cases, showing a clear advantage over the original names. We provide more details of our human study in the supplements.
4.2 Upgrading benchmarks with renovated names
Setup. We demonstrate that renovated names can be used to upgrade existing benchmarks on two popular open-vocabulary benchmarks, ADE20K (ADE-150) and Cityscapes. Thanks to our renovated names, our upgraded benchmarks contain 4 to 5 times more classes than the original ones. Besides being more fine-grained, our benchmarks are also more realistic, aligning more closely with how humans use names to prompt models. As we show later, these renovated names provide valuable extra information to the existing benchmarks.
To ensure high quality names for segments in the validation sets, we involve five human annotators in a rigorous verification process. Specifically, we present human annotators with the top 3 name suggestions along with their confidence scores from our renaming model. We then ask them to select the best-matching names among them. For the few segments that do not match with any of the top 3 names, the annotators will check the whole candidate name list of the corresponding groundtruth class. In case of disagreements over the name selection, all the annotators discuss and decide the best choices. We will release these annotations upon publication.
Along with the benchmarks, we incorporate new evaluation metrics [58] that account for the semantic similarity between names, rather than the standard evaluation metrics of panoptic quality (PQ), Average Precision (AP), and mean Intersection-over-Union (mIoU) which penalize the classification errors in a “0/1” manner. For example, under standard metrics, predicting an “SUV” as a “car” will be penalized as much as erroneously labeling it as a “person”. This significantly deviates from the nuanced understanding humans apply when categorizing objects. To better align with humans, we adopt the recently proposed “open” metrics [58] for our upgraded benchmarks. In particular, for a groundtruth mask (pixel) $g_{i}$ with its label $c_{i}$ and the predicted mask (pixel) $d_{j}$ with its predicted label $c_{j}$, we compute soft $TP_{c_{i}}=S_{c_{i},c_{j}}$, soft $FP_{c_{j}}=1-S_{c_{i},c_{j}}$, and soft $FN_{c_{i}}=1-S_{c_{i},c_{j}}$, where $S_{c_{i},c_{j}}$ is the semantic similarity between $c_{i}$ and $c_{j}$. These soft TP, FP, and FN are then used to compute metrics PQ, AP, and mIoU in a standard way. With the “open” metrics, the evaluation of open-vocabulary segmentation would be more aligned with human perception, where closely related names incur lower penalties compared to those that are significantly different.
ADE20K [57] Cityscapes [7]
Benchmark Test Name Metric #classes PQ AP mIoU #classes PQ AP mIoU
Original Original Standard 150 24.30 16.79 32.14 19 39.42 22.76 53.08
Open [58] 150 27.04 16.52 36.59 19 42.57 26.54 62.69
OpenSeg Standard 150 25.35 17.30 33.06 19 43.68 26.93 56.15
Open [58] 150 29.03 17.52 41.03 19 45.92 27.09 60.21
Upgraded RENOVATE Standard 578 7.07 5.42 5.83 108 8.62 8.63 7.84
Open [58] 578 25.37 12.48 16.73 108 27.17 24.92 32.25
Table 2: Open-vocabulary segmentation on upgraded benchmarks. Our upgraded benchmarks with RENOVATE names contain more fine-grained annotations and are more challenging. Here we use pre-trained FC-CLIP model for evaluation.
In the original benchmarks, the evaluation metrics are computed on the original class division (150 classes for ADE20K and 19 classes for Cityscapes). In our upgraded benchmarks, the text prompt is the set of all the unique names in the datasets (578 classes for ADE20K and 108 classes for Cityscapes). When computing evaluation metrics, we also use this set of names as class division. Thus, the upgraded benchmarks contain much more classes than the original benchmarks. Since our names match better with the visual segments, the upgraded benchmarks are more realistically reflecting how human users may prompt the model for segmentation.
Results. In Tab. 2, we evaluate FC-CLIP model [52] on the upgraded benchmarks of ADE20K and Cityscapes and compare with the evaluation results on the original benchmarks. We also experiment with two other models, ODISE [48] and MasQCLIP [51], which lead to similar findings. See results in supplements.
We first observe that our benchmark is more challenging, as evidenced by considerably lower performance scores. This is because our benchmark requires the model to make classification on much more classes. This means predictions are closer to the decision boundaries and are thus easier to confuse between classes, leading to higher classification errors.
We also note that the discrepancy between standard and open metrics is much greater in our upgraded benchmarks. While standard benchmarks show minimal difference due to their class names being more distinct from each other, our benchmarks highlight a significant underestimation of the model’s open-vocabulary capability since the renovated names can be very similar. Open metrics address this by factoring in semantic similarity, providing a more realistic performance measure on our more fine-grained benchmarks.
Figure 6: Per-category IoU on the upgraded ADE20K with 578 classes. We highlight the top/bottom-5 renovated names and some selected names from the original “person” class. Our upgraded benchmarks allow us to conduct more fine-grained analysis on open-vocabulary model abilities and biases.
Finally, we demonstrate in Fig. 6 that our upgraded benchmarks enable a more detailed examination of models’ open-vocabulary capabilities and inherent biases. In particular, we find that the pre-trained model excels in identifying categories like “sky”, “girl”, and “sports car” with high mIoU scores, while struggling with “bollard”, “ceiling light”, and “tray”, reflecting the trend that large, well-defined, and common objects tend to be segmented more accurately than those that are small, deformable, and rare. Note that most of these names do not exist in the original benchmarks. In addition, our analysis reveals model biases, such as the disparity in mIoU scores between “man” and “woman”, “girl” and “boy”, suggesting an imbalance in the training dataset. Such fine-grained information is pivotal for improving models and informing dataset curation, demonstrating the substantial advantages of our upgraded benchmarks.

### v2

@@ Abstract
Names are essential to both human cognition and vision-language models. Open-vocabulary models utilize class names as text prompts to generalize to categories unseen during training. However, the precision of these names is often overlooked in existing datasets. In this paper, we address this underexplored problem by presenting a framework for “renovating” names in open-vocabulary segmentation benchmarks (RENOVATE). Our framework features a renaming model that enhances the quality of names for each visual segment. Through experiments, we demonstrate that our renovated names help train stronger open-vocabulary models with up to 15% relative improvement and significantly enhance training efficiency with improved data quality. We also show that our renovated names improve evaluation by better measuring misclassification and enabling fine-grained model analysis. We will provide our code and relabelings for several popular segmentation datasets (MS COCO, ADE20K, Cityscapes) to the research community.
1 Introduction
@@
Figure 1: Problems of names in current segmentation benchmarks. We demonstrate examples from well-known datasets: MS COCO [13], ADE20K [14], and Cityscapes [15]. Our renovated names are visually more aligned and help models to generalize better.
In this work, we focus on renovating the names for open-vocabulary segmentation benchmarks, as their dense annotations pose a greater challenge and offer broader applicability than other recognition tasks. The importance of names in open-vocabulary segmentation is often overlooked, with one exception, OpenSeg [5], which manually inspected and modified the class names of several segmentation benchmarks, resulting in considerable performance gains of their proposed model. While their work demonstrated the importance of choosing good names, a manual approach is subjective and difficult to scale. To date, there exists no systematic study on how to rename benchmarks in a scalable and principled way. Our work is the first attempt to tackle this challenge.
In this work, we introduce a scalable, simple, yet general method for renaming segmentation benchmarks that outperforms manual label efforts like OpenSeg [4]. Our approach leverages foundation models to automate the renaming process, significantly reducing the manual labor traditionally required. Towards this goal, our method can be divided into two steps. First, we narrow down the name search space from the whole language space to a curated list of candidate names. This list is generated by leveraging the original class names with contextually relevant nouns extracted from visual contents via an image captioning model, thereby streamlining the search for names. Next, we employ a specially trained renaming model to identify and select the best-matching candidate name for every ground-truth segmentation mask. In this way, we match an individual name for each instance without any extra human annotations.
To demonstrate the value of RENOVATE names, we explore two practical applications. We first use our renovated names for training open-vocabulary models and show that they substantially enhance generalization performance across various benchmarks by up to 15%, indicating a more accurate alignment with visual segments. Our new names also significantly increase training efficiency thanks to their enriched textual information. Second, we apply our renovated names to improve evaluation of open-vocabulary segmentation. Specifically, we reveal that current pre-trained models are making many “benign” misclassifications, i.e., misclassified names that are semantically close to the ground-truth names. RENOVATE names enable more fine-grained model performance analysis, providing valuable information for further model enhancement and dataset curation.
@@
  • We point out the naming problem in existing open-vocabulary segmentation benchmarks.
  • We propose a scalable, simple, and general framework for automatic dataset renaming.
  • We demonstrate that our RENOVATE names help to train models with stronger open-vocabulary capabilities and refine evaluation benchmarks, providing valuable insights for future research.
@@ 3.1 Generating candidate names
We use GPT-4 [29] for creating a pool of candidate names. A naive solution is to prompt GPT-4 with the original name and ask it to generate synonyms and hierarchical concepts. However, since the original names are often too general and ambiguous, GPT-4 does not have sufficient knowledge to generate high-quality candidates. Therefore, we propose to exploit the visual contents for generating some “context names” (Table 1) that assist GPT-4 in comprehending the category’s meaning prior to generating candidate names.
@@
Specifically, we instruct GPT-4 to generate two types of candidate names: (1) Synonyms or subcategories, which can expand the lexical scope of the overly general class name. For instance, the class name “grass” could yield variations like “lawn” or “turf”. For this type, synonyms and subcategories are considered together, acknowledging their often interchangeable nature. (2) Providing short, distinct contexts to polysemous names. As an example, the ambiguous name “fan” could be elaborated into “ceiling fan” or “floor fan”, offering clearer specificity. As shown in Table 1, the 5-10 candidate names generated by GPT-4 are more precise than the original names, reflecting the context information encoded by context names. We provide detailed GPT-4 prompts in the supplement.
@@ 3.2 Training for candidate name selection
Among the generated candidates, we will only keep those that are better aligned with the visual contents. Particularly, we train a model that can assess the alignment between each name and segment, using segmentation as the proxy task. Conceptually, a candidate name that well describes the segment in the image should help the model recover its ground-truth mask and make a correct classification. Therefore, the model should allow vision-language interaction, and the supervision signal should encourage to use textual information for mask prediction and classification.
Our renaming model is illustrated in Fig. 2. Its vision part starts from a CLIP-based vision backbone followed by a pixel decoder that gradually upscales the backbone features to generate per-pixel embeddings $F_{\text{pix}}\in\mathbb{R}^{H\times W\times C}$. The candidate names of the categories present in the image are processed by the CLIP text encoder. The interaction of textual and visual information takes place at the transformer decoder. The initial queries $X^{(0)}$ are from the text embeddings, and then updated by the transformer decoder consisting of $L$ blocks using the multi-scale visual features from the pixel decoder. As output of each block, the queries $X^{(l)}$ generate intermediate mask and class predictions:
@@
where * refers to pixel-wise multiplication. The final predictions are made by the queries $X^{(L)}$. Our pixel decoder, mask prediction and classification follow the design of Mask2Former [31], which handles panoptic, instance, and semantic segmentation in a unified manner. The key differences lie in 1) the transformer decoder design, and 2) the training strategy, which we detail subsequently.
Transformer Decoder. Our transformer decoder uses text embeddings as input queries, unlike most open-vocabulary segmentation models that employ text embeddings solely as weights of the Linear layer in Eq. 1 for classification. This key change fosters an earlier integration of text and visual features at varying scales, essential for leveraging text information in segmentation tasks. As shown in Fig. 2, the text embedding of each candidate name from the segment’s class “bicycle” feeds into the transformer decoder, generating their corresponding mask and class predictions.
@@
Training Strategy. With the CLIP backbone kept frozen, we train the transformer decoder together with the pixel decoder. To recover the ground-truth mask $M_{i}$ and class $c_{i}$, the transformer decoder makes multiple mask and class predictions with candidate names from the class $c_{i}$. The prediction with the highest Intersection over Union (IoU) with $M_{i}$ is selected for loss computation:
@@ 3.3 Obtaining renovated names
As illustrated in Fig. 3(a), to obtain the renovated names, we run the trained renaming model to associate each ground-truth mask with the candidate name. For each segment, our renaming model first ranks the names by the IoU scores between the predicted masks and the ground-truth mask, then selects the top-scoring name as the renovated name. While the IoU ranking scores are not probabilities, they are valuable to understand the relative and absolute model confidence in the names, i.e., how the chosen name stands against the other candidates and whether the selection is indeed confident. In the experiments (Section 5.3), we make use of this information to speed up human verification. We show some examples of renovated names in Fig. 4 and more in the supplements.
@@
Our renovated names are readily usable for analyzing the dataset distribution and biases in a more fine-grained manner, as exemplified in Fig. 3(b). Note that our renovated names in each original class are not necessarily mutually exclusive because different visual segments may be matched to names at different hierarchies. For example, in Fig. 4, a person is named as “engineer” (second row), likely due to the clothing, while the other person is named as “woman” (third row), possibly from the physical appearance. But “engineer” and “woman” are not mutually exclusive names. An interesting future extension would be to further study the hierarchies of different names in a dataset.
4 Applications of RENOVATE Names
We further demonstrate the value of the renovated names with two application tasks: training stronger open-vocabulary segmentation models and improving existing evaluation benchmarks. The improvements in both tasks indicate the quality of RENOVATE names.
4.1 Training with RENOVATE names
RENOVATE names offer precise and diverse text information. We exploit this to improve model training and achieve better generalization and data efficiency. Following prior works [24, 25, 33], we replace original ground-truth names with RENOVATE names to compute text embeddings. In addition, we employ the “negative sampling” technique, commonly used in natural language processing to enhance training with a large number of classes [34, 35, 36, 37]. For each segment, we first randomly select $C-1$ negative classes and choose one RENOVATE name from each. These names are then concatenated with the ground-truth RENOVATE name to form the text prompt. This effectively constructs text embeddings of length $C$ with semantically distinct names, improving training efficiency and generalization. For computing the classification loss, we use the cross entropy function on a per-segment basis, i.e., for an image with segments $\{m_{i}\}_{i=1}^{N}$, $\mathcal{L}_{\text{class}}=\sum_{i=1}^{N}\text{CE}(c_{i},\text{softmax}(v_{i}T_{i}^{T})),$ where $c_{i}$ is the ground-truth label, $v_{i}\in\mathbb{R}^{d}$ is the predicted visual embeddings and $T_{i}\in\mathbb{R}^{C\times d}$ is the constructed text embeddings through negative sampling. Throughout our experiments, we use negative sampling in our model training unless specified.
4.2 Improving evaluation with RENOVATE names
RENOVATE names also enable more fine-grained analysis of open-vocabulary models. For instance, for a misclassification from “car” to “van”, we may distinguish between “SUV” to “minivan” versus “sedan” to “delivery van”, which isn’t possible with class-level annotations in current datasets. However, even with RENOVATE names, standard segmentation metrics obscure our desired fine-grained performance insights since they penalize all such misclassification cases equally. To this end, we adopt “open” evaluation metrics [38] that consider semantic similarity between names. Specifically, for a groundtruth mask pixel $g_{i}$ with label $c_{i}$ and a predicted mask pixel $d_{j}$ with predicted label $c_{j}$, we compute soft $TP_{c_{i}}=S_{c_{i},c_{j}}$, soft $FP_{c_{j}}=1-S_{c_{i},c_{j}}$, and soft $FN_{c_{i}}=1-S_{c_{i},c_{j}}$, where $S_{c_{i},c_{j}}$ is the semantic similarity between $c_{i}$ and $c_{j}$. These soft metrics are then used to compute panoptic quality (PQ), Average Precision (AP), and mean IoU (mIoU) in a standard way. In Section 5.3, we demonstrate how open metrics effectively reveal that there are indeed many “benign” misclassification scenarios using RENOVATE names.
While our renaming process is automated, it is essential to ensure exceptionally high-quality names for evaluation. Therefore, we introduce a rigorous verification process involving five human annotators who review the top three name suggestions and their confidence scores from our model to select the best match. If none of the top three names are suitable, verifiers examine the entire candidate list. In case of disagreements, verifiers discuss and decide collectively. Note that human verification is used only for evaluation sets, while automatically generated names are used for the much larger training sets. Future work could explore using multiple vision-language foundation models to automate this verification process as evaluation sets grow [39, 40, 41].
@@ 5 Experiments
5.1 Obtaining renovated names
@@
Setup. We renovate three panoptic segmentation datasets respectively: MS COCO [13], ADE20K [14], and Cityscapes [15]. To generate context names, our default image captioning model is CaSED [30], which retrieves captions by matching vision embeddings with text embeddings of captions from a large-scale PMD dataset [42]. We train a renaming model for 60k iterations with a batch size of 16 on the training set, then generate RENOVATE names for the entire dataset.
COCO ADE CS
# Original classes 133 150 19
# Segments/Class 11,274 2,009 4,964
# RENOVATE names 741 578 108
# Segments/Name 1,871 521 873
Table 2: Statistics of renovated datasets.
Results. Table 2 summarizes the statistics of RENOVATE names. Compared to the original class names, each dataset has approximately 5-6 times more distinct RENOVATE names, indicating that they provide more diverse and fine-grained semantic annotations to the visual segments. We show some visual examples in Fig. 4 and more in Section E.1. To further validate the quality of the names, we conduct a human preference study, see Section B.1. We also conduct an ablation analysis on the components of the renaming pipeline, see Section B.2.